OpenStudy (anonymous):

Solve, using the substitution method. y + 2x = 7 14 – 4x = 2y A.The solution is (1, 5) B. The solution c. There are an infinite number of solutions is (21, 0) D.There is no solution. C.

OpenStudy (anonymous):

Using substitution: solving the first equation for y, y = 7 - 2x substituting it into the second equation: 14 - 4x = 2(7 -2x) 14 - 4x = 14 - 4x ... they are the same equation so there are infinitely many solutions